A manufacturer’s job cost sheet reports direct materials of $1,200 and direct labor of $250 for...

  1. A manufacturer’s job cost sheet reports direct materials of $1,200 and direct labor of $250 for printing 200 T-shirts for a bikers’ reunion. Estimated overhead is computed as 140% of direct labor costs.
    1. What is the estimated overhead cost for this job?
    2. What is the total cost per T-shirt for this job?
    3. What journal entry does the manufacturer make upon completion of this job to transfer costs from work in process to finished goods?
  1. A manufacturing company estimates it will incur $240,000 of overhead costs in the next year. The company applies overhead using machine hours and estimates it will use 1,600 machine hours in the next year. During the month of June, the company used 80 machine hours on job 1 and 70 machine hours on Job 2.
    1. Compute the predetermined overhead rate to be used to apply overhead during the year.
    2. Determine how much overhead should be applied to Job 1 and to Job 2 for June
    3. Prepare the journal entry to record overhead applied for June.
Answer #1

1) Estimated overhead cost for this Job = 250*1.4 = 350

Total cost per T-shirt = (1200+250+350)/200 = 9

2) Journal entry

No General Journal Debit Credit
Finished goods 1800
Work in process 1800

3) Overhead rate = 240000/1600 = 150 per MH

Overhead applied :

Job 1 = 80*150 = 12000

Job 2 = 70*150 = 10500

Journal entry

No General Journal Debit Credit
Work in process (12000+10500) 22500
Manufacturing overhead 22500
(To overhead applied)